# Channel‑Selected Stratified Nested Cross‑Validation for Clinically Relevant EEG‑Based Parkinson’s Disease Detection

### v1

@@ Abstract
The early detection of Parkinson’s disease remains a critical challenge in clinical neuroscience, with electroencephalography offering a non‑invasive and scalable pathway toward population‑level screening. While machine learning has shown promise in this domain, many reported results suffer from methodological flaws, most notably patient‑level data leakage, inflating performance estimates and limiting clinical translation. To address these modeling pitfalls, we propose a unified evaluation framework grounded in nested cross‑validation and incorporating three complementary safeguards: (i) patient‑level stratification to eliminate subject overlap and ensure unbiased generalization, (ii) multi‑layered windowing to harmonize heterogeneous EEG recordings while preserving temporal dynamics, and (iii) inner‑loop channel selection to enable principled feature reduction without information leakage. Applied across three independent datasets with a heterogeneous number of channels, a convolutional neural network trained under this framework achieved 80.6% accuracy and demonstrated state‑of‑the‑art performance under held‑out population block testing, comparable to other methods in the literature. This performance underscores the necessity of nested cross‑validation as a safeguard against bias and as a principled means of selecting the most relevant information for patient‑level decisions, providing a reproducible foundation that can extend to other biomedical signal analysis domains.
@@ I Introduction
Parkinson’s disease (PD) is the second most common neurodegenerative disorder, affecting over 10 million people worldwide and imposing a growing socioeconomic burden [1]. Its progressive course—marked by motor dysfunction, cognitive decline, and reduced quality of life—makes early detection and intervention a critical priority  [2, 3]. Yet current clinical diagnosis is largely based on overt motor symptoms, which typically emerge only after substantial dopaminergic neuron loss [2, 3]. This delay highlights the importance of identifying prodromal biomarkers, such as REM sleep behavior disorder, olfactory impairment, and subtle motor irregularities, which often precede clinical diagnosis by years [2]. Detecting early PD could enable interventions when therapies are effective [2].
Advances in biomedical signal acquisition have created new opportunities for non‑invasive monitoring of these prodromal and progressive symptoms [4]. Among available modalities, electroencephalography (EEG) is particularly promising: it is portable, relatively inexpensive, and offers millisecond‑level temporal resolution of neural dynamics [5]. Compared to imaging techniques such as MRI or PET, which provide high spatial resolution but are costly and less accessible, EEG enables scalable, repeated assessments that are well‑suited for longitudinal monitoring [5, 6]. This combination of accessibility and temporal precision positions EEG as a strong candidate for early PD biomarker discovery.
@@
Figure 1: The nested cross‑validation architecture addresses the challenge of high‑dimensional EEG data by starting with all channels and systematically reducing to the most informative. The inner loop selects channels contributing most to inference, while the outer loop validates this choice without bias. This principled “many‑to‑best” reduction is dataset‑, channel‑, and model‑agnostic, yielding results that are more efficient, generalizable, and interpretable.
@@
In parallel, machine learning (ML) and deep learning (DL) methods show strong potential for extracting subtle biomarkers from high‑dimensional biomedical data [7]. A wide range of neural architectures have been explored for biosignal analysis: recurrent neural networks (RNNs) and long short‑term memory (LSTM) models capture temporal dependencies in sequential data [8]; transformer‑based models leverage self‑attention to model long‑range interactions across time and frequency; and graph neural networks (GNNs) exploit the spatial topology of EEG electrode layouts to learn structured representations [9]. Despite these advances, convolutional neural networks (CNNs) remain a principled starting point for EEG analysis. When applied to short‑time Fourier transform (STFT) spectrograms, CNNs exploit partial shift‑invariance that complements the local stationarity of the STFT, enabling robust feature extraction without handcrafted preprocessing [10, 11, 12]. This synergy has consistently delivered state‑of‑the‑art performance across diverse signal domains, including PD detection, where CNN‑based or hybrid pipelines frequently report accuracies exceeding 95% [13]. Moreover, CNN architectures are naturally extensible to multimodal fusion: EEG inputs can be paired with bio-signal sensor data through parallel branches converging in shared layers, allowing integrated biomarker discovery without redesigning the core model [14].
Many reported successes in EEG‑based PD detection are undermined by methodological flaws, most notably data leakage and dataset heterogeneity. Leakage arises in two primary forms: temporal leakage, where overlapping windows from the same recording are split across training and test sets, and subject leakage, where data from the same patient appears in both folds. In both cases, models risk memorizing idiosyncratic traits rather than learning disease‑relevant features, inflating performance and eroding clinical validity [15]. Surveys of PD classification studies further highlight that EEG research often employs ambiguous patient‑level stratification, if it is applied at all [16]. Population block testing offers a partial safeguard, yet its scope is limited, as models may still overfit to the unique characteristics of a single cohort. These concerns are compounded by the heterogeneity of public EEG datasets: channel counts range from high‑density 64‑channel caps to sparse consumer devices, and acquisition hardware differs in sampling rate, electrode placement, and preprocessing pipelines. Methods tuned to one configuration often fail to transfer across others, creating a significant reproducibility barrier. Collectively, these challenges underscore the need for architectures that prevent leakage, generalize across populations, and robustly reduce high‑dimensional signals to the most informative subset, independent of dataset or recording setup.
@@
To address these challenges, we propose a nested cross-validation framework for robust EEG analysis (Figure 1). The outer loop enforces patient-level stratification to eliminate leakage, while the inner loop selects the most informative channels, reducing dimensionality without sacrificing clinically relevant signals. A multi-layered windowing strategy integrates recordings of varying length, sampling rate, and hardware, enabling consistent evaluation across heterogeneous datasets. By design, the framework is dataset, channel, and model-agnostic, positioning it as a blueprint for reproducible EEG-based biomarker discovery. Together, these components yield generalizable performance estimates and clinically meaningful early PD detection results.
Our contributions are threefold:
  1. 1.
    A model‑agnostic nested cross‑validation framework with stratification, windowing, and inner‑loop channel selection to prevent leakage, harmonize datasets, and reduce dimensionality.
  2. 2.
    Empirical evidence that patient‑level leakage and narrow cohort blocking inflate EEG‑based PD detection performance and undermine clinical validity.
  3. 3.
    Demonstration that learned representations align with canonical EEG frequency bands, yielding interpretable models that achieve 0.806 accuracy across datasets.
@@
Looking forward, this framework provides a foundation for accelerating clinical translation: it enables reproducible biomarker validation across heterogeneous cohorts, supports integration with multimodal data streams such as EMG and wearable sensors, and lays the groundwork for scalable digital health pipelines capable of monitoring neurodegenerative disease progression in real-world settings.
@@ II Related Work and Methodological Pitfalls
A growing body of research has applied ML techniques to the early detection of PD using EEG, EMG, and other biosignal modalities. Several studies have reported high classification accuracies, often exceeding 85%, suggesting strong potential for clinical translation [17, 18]. However, closer inspection reveals that many of these works rely on evaluation protocols that do not stratify data at the patient level. Instead, temporal windows from the same subject are distributed across both training and test sets, creating a form of data leakage. This allows models to exploit subject‑specific idiosyncrasies rather than learning generalizable disease markers, thereby inflating performance estimates [15, 19, 20, 21].
Beyond this central issue, several additional pitfalls recur across the literature. The small sample sizes typical of PD datasets magnify overfitting risks, as even a few misallocated subjects can distort reported outcomes [20]. Reported performance is also often reduced to accuracy, which obscures clinically relevant trade‑offs; complementary metrics such as sensitivity, specificity, ROC‑AUC, and calibration curves provide a more reliable assessment [22, 23]. Preprocessing choices—including channel selection, filtering, and artifact removal—introduce further variability and can bias results when hyperparameters are tuned on overlapping data [19, 24]. Although nested cross‑validation is widely used in other domains for ensembling [19] and hyperparameter optimization [25], it can be applied in PD detection to quantify variability across patient‑level hold‑out sets and channel configurations. In heterogeneous settings with rare events, neglecting such stability checks risks overconfident claims of deployment readiness [21]. Finally, no existing PD studies attempt non‑harmonized cross‑dataset validation (e.g., Iowa vs. San Diego), leaving open whether models trained on one population can generalize to others with different acquisition protocols or demographics.
DL approaches have also been explored for PD detection, but many inherit the same evaluation flaws. For instance, [26] applied convolutional neural networks to EEG recordings of patients and controls with ambiguous enforcement of subject‑independent splits, leaving the reported performance difficult to interpret. Likewise, [13] presented hybrid convolutional models with seemingly strong results, yet overlapping patient data across folds assuredly inflated accuracy estimates. Comparable practices appear in other studies [16, 27, 28, 21], where non‑stratified cross‑validation or single‑block testing at the population level introduce systematic bias. Collectively, these works illustrate how even sophisticated methods can yield misleading outcomes when evaluation fails to guard against leakage and overfitting.
Figure 2: Suggested Workflow for EEG Channel Selection Modeling: Subjects are represented by raw EEG time‑series, which are transformed into spectrograms via short‑time Fourier transform. An inner cross‑validation loop trains a model to estimate per‑channel performance, retaining only the most informative electrodes. These selected channels are reintegrated into the outer training and validation framework, where the final model is trained. Predictions are then aggregated into a single patient‑level probability, providing a concise outcome measure.
A more recent example is the GAN‑Enhanced Generalizable Model for EEG‑Based Detection of Parkinson’s Disease (GEPD) [29], which reported roughly 85% accuracy using generative augmentation. While advancing cross‑dataset generalization, its reliance on a single 90/10 split and train‑on‑Iowa/test‑on‑UNM configuration—without patient‑level aggregation—raises concerns. Similar protocols, including the reverse train/test setup [16, 21], risk tuning models to population‑specific artifacts, producing unstable and inflated performance estimates. Furthermore, GEPD discarding non‑matching EEG channels across datasets limits clinical applicability. More broadly, generative augmentation itself can amplify inherent biases if synthetic samples reproduce population‑specific artifacts rather than disease‑relevant features, underscoring the need for careful validation [30].
In contrast, our design emphasizes rigorous evaluation by integrating three public datasets without discarding viable channels, applying five‑fold stratified cross‑validation with patient grouping, and reporting outcomes at the patient level. These protocols mitigate overfitting to dataset artifacts and yield more reliable estimates of generalization, a critical requirement for biomedical applications. Our approach also advances interpretability, where prior studies treat neural networks as black boxes, we provide band‑level explanations of EEG features, Illuminating which spectral components drive predictions. Thus, methodological robustness and transparent feature attribution demonstrates that competitive performance can be achieved without data leakage or singular validations, while also enhancing trust and clinical relevance.
@@ III-A Preprocessing
All preprocessing steps are intended to normalize EEG signals within target frequency bands for our nested cross-validation framework. Let $\mathbf{s}_{m}(t)$ denote the raw EEG signal for patient $p_{m}$, sampled at original frequency $f_{\text{orig}}$. Each signal was resampled to a common rate of $f_{s}=64~\text{Hz}$ to preserve oscillatory activity in the Beta (13133030 Hz) and Theta (4488 Hz) bands. The resampled signal $\tilde{\mathbf{s}}_{m}(t)$ was standardized to zero mean and unit variance, $\mathbf{z}_{m}(t)=\frac{\tilde{\mathbf{s}}_{m}(t)-\mu_{m}}{\sigma_{m}},$ where $\mu_{m}$ and $\sigma_{m}$ are the mean and standard deviation of $\tilde{\mathbf{s}}_{m}(t)$. Each standardized trace was then normalized to unit amplitude, $\hat{\mathbf{s}}_{m}(t)=\frac{\mathbf{z}_{m}(t)}{\max_{t}|\mathbf{z}_{m}(t)|}.$ For each subject $p_{m}$, the aligned EEG arrays $\hat{\mathbf{s}}_{m}(t)$ were stored in a DataFrame with metadata (ID, session, diagnosis, file name). Thus, ensuring reproducibility, subject-level traceability, and partitioning within outer cross-validation folds, preventing leakage from the outset.
To ensure comparability across datasets, we incorporated a channel harmonization step. Let $\mathcal{C}=\{c_{1},c_{2},\dots,c_{C}\}$ denote the full set of electrodes, where $|C|$ corresponds to the number of channels in the Iowa dataset. Channels were first partitioned into anatomical regions, $\mathcal{C}=\mathcal{C}_{\text{frontal}}\cup\mathcal{C}_{\text{fronto‑central}}\cup\mathcal{C}_{\text{central}}\cup\cdots\cup\mathcal{C}_{\text{occipital}},$ with each subset $\mathcal{C}_{r}$ preserving spatial contiguity. Within each region, electrodes were ordered according to standard EEG conventions:
@@
After ordering, each channel was assigned a unique integer index $\pi:\mathcal{C}\rightarrow\{0,1,\dots,C-1\},$ such that $\pi(c)$ reflects its position in the harmonized left‑to‑right, anterior‑to‑posterior sequence. This mapping provided a consistent spatial template across datasets, enabling reproducible alignment of electrodes to model inputs and ensuring that subsequent channel selection operated on harmonized, regionally structured arrays. To accommodate datasets with incomplete coverage, any channel $c\in\mathcal{C}$ absent from a given recording was zero-padded at its corresponding index $\pi(c)$. This ensured that all input arrays conformed to the full Iowa-based template, preserving spatial structure while maintaining compatibility across heterogeneous datasets.
III-B Cross-Validation Schema
As illustrated in Figure 2, we adopt a patient‑stratified nested cross‑validation design where the outer loop estimates generalization to unseen patients and the inner loop supports model selection and channel pruning. Both loops use StratifiedGroupKFold with patient identifiers from our DataFrame’s metadata as grouping variables, ensuring all samples $\mathcal{W}_{m}$ from a patient $p_{m}$ are confined to either training or testing folds. This structure prevents patient and temporal leakage while preserving class balance, whereby nesting hyperparameter tuning and channel selection within the inner loop maintains unbiased outer evaluations.
III-C Windowing and Feature Extraction
To define the data segments used in the cross‑validation schema, we employed a windowing strategy tailored to split the ‘Full EEG Signal Stack’ of each patient into channel separated spectrograms, as depicted in Figure 2, which vary in length and recording conditions. With sampling rate $f_{s}=64~\text{Hz}$, the outer window length was set to $L_{w}=16{,}384~\text{samples},$ corresponding to $\frac{16{,}384}{f_{s}}\approx 256~\text{s}.$ Shorter signals were zero‑padded, and longer signals segmented into overlapping windows with hop size $H_{w}=\frac{L_{w}}{4}=4{,}096~\text{samples}.$ Within each window, we computed the short‑time Fourier transform (STFT). For a discrete‑time signal $\mathbf{w}[n]\in\mathbb{R}^{T}$, the STFT is
@@ III-C Windowing and Feature Extraction
where $L$ is the analysis window length, $b$ the hop size, $\mathbf{x}[k]$ the window function, $f$ the frequency bin index, and $i$ the time frame index. This yields a complex‑valued matrix of size $N\times L$, with $N=\left\lfloor\frac{T}{b}\right\rfloor,\quad\Delta f=\frac{f_{s}}{L}.$ We used $n_{\text{fft}}=256$ and $h=64$ samples, producing spectrograms with $F=\frac{n_{\text{fft}}}{2}=128\quad\text{frequency bins},\qquad T=\frac{L_{w}}{h}=\frac{16{,}384}{64}=256\quad\text{time bins}.$ Each EEG channel was thus represented as $(F\times T\times 1)=(128\times 256\times 1),$normalized and converted to decibel scale. This procedure expands each patient’s EEG into multiple channel separated spectrograms, increasing training sample size and enabling finer‑grained decisions while maintaining stable computational demands.
III-D Channel Selection
As highlighted in the ‘Inner Loop Channel Selection’ stage of Figure 2, channel selection is embedded within the nested cross‑validation schema to address the high dimensionality of EEG recordings, where $C\in\{32,64\}$ active electrodes capture distinct neural activity. For each outer fold’s training set, spectrograms from all active channels $\{c_{1},c_{2},\dots,c_{C}\}$ are passed through the CNN, and per‑channel accuracies $a_{c,f}$ are recorded on each inner‑fold validation subset $f\in\{1,\dots,F\}$. These fold‑wise accuracies are then accumulated and averaged to obtain a stability‑adjusted score for each channel where only the active channels for each patient are accumulated into their respective average,

### v2

@@ Abstract
The early detection of Parkinson’s disease remains a critical challenge in clinical neuroscience, with electroencephalography (EEG) offering a non‑invasive and scalable pathway toward population‑level screening. While machine learning has shown promise in this domain, many reported results suffer from methodological flaws, most notably patient‑level data leakage, inflating performance estimates and limiting clinical translation. To address these modeling pitfalls, we propose a unified evaluation framework grounded in nested cross‑validation and incorporating three complementary safeguards: (i) patient‑level stratification to eliminate subject overlap and ensure unbiased generalization, (ii) multi‑layered windowing to harmonize heterogeneous EEG recordings while preserving temporal dynamics, and (iii) inner‑loop channel selection to enable principled feature reduction without information leakage. Applied across three independent datasets with a heterogeneous number of channels, a convolutional neural network trained under this framework achieved 80.6% accuracy and demonstrated state‑of‑the‑art performance under held‑out population block testing, comparable to other methods in the literature. This performance underscores the necessity of nested cross‑validation as a safeguard against bias and as a principled means of selecting the most relevant information for patient‑level decisions, providing a reproducible foundation that can extend to other biomedical signal analysis domains.
@@ I Introduction
Parkinson’s disease (PD) is the second most common neurodegenerative disorder, affecting over 10 million people worldwide and imposing a substantial socioeconomic burden, yet clinical diagnosis typically relies on overt motor symptoms that emerge only after significant neuronal loss [1, 2, 3]. This delay underscores the need to identify prodromal biomarkers—such as REM sleep behavior disorder and subtle motor changes—that may precede diagnosis by years and enable earlier intervention [2]. Recent advances in biomedical signal acquisition have made such early monitoring increasingly feasible, with EEG emerging as a particularly promising modality for non‑invasive assessment of prodromal and progressive PD symptoms due to its portability, low cost, and high temporal resolution [4, 5]. Unlike MRI or PET, EEG enables scalable, repeated assessments, supporting longitudinal monitoring and biomarker discovery [5, 6].
@@
Figure 1: The nested cross‑validation architecture is applied to high‑dimensional EEG by training models on all channels, while inner‑loop validation ranks channel relevance and defines the optimal test‑time architecture. The outer loop evaluates this selected configuration, enabling a dataset‑, channel‑, and model‑agnostic “many‑to‑best” reduction that supports efficient, generalizable, and interpretable inference.
@@
In parallel, machine learning (ML) and deep learning (DL) methods show strong potential for extracting subtle biomarkers from high‑dimensional biomedical data [7]. A wide range of neural architectures have been explored for biosignal analysis, including temporal recurrent models, self‑attention based transformer models, and graph neural networks that exploit EEG electrode layouts [8, 9]. Despite these advances, convolutional neural networks (CNNs) remain a principled starting point for EEG analysis. When applied to short‑time Fourier transform (STFT) spectrograms, CNNs exploit partial shift‑invariance aligned with the local stationarity of the STFT [10, 11, 12]. This synergy has consistently delivered state‑of‑the‑art performance across diverse signal domains, including PD detection, where CNN‑based or hybrid pipelines frequently report accuracies exceeding 95% [13]. Moreover, CNN architectures are extensible to multimodal fusion, allowing EEG to be integrated with complementary biosignals through parallel branches [14].
Many reported successes in EEG‑based PD detection are undermined by methodological issues, particularly data leakage and dataset heterogeneity. Leakage from overlapping temporal windows or subject overlap allows models to exploit patient‑specific idiosyncrasies rather than disease‑relevant features [15], and population‑level splits provide only partial protection against cohort‑specific overfitting [16]. Combined with heterogeneity in EEG channel counts, hardware, and preprocessing, these factors limit reproducibility and cross‑dataset transferability, motivating evaluation frameworks that explicitly prevent leakage and assess generalization across diverse populations.
@@
To address these challenges, we propose a nested cross‑validation framework for robust EEG analysis (Figure 1) that enforces patient‑level stratification in the outer loop to eliminate leakage, performs channel selection in the inner loop to reduce dimensionality without discarding clinically relevant signals, and integrates a multi‑layered windowing strategy to accommodate recordings with varying length, sampling rate, and hardware. By design, the framework is dataset‑, channel‑, and model‑agnostic, positioning it as a reproducible blueprint for EEG‑based biomarker discovery. Accordingly, our contributions are threefold: (i) a model‑agnostic, stratified nested cross‑validation framework with integrated channel selection to prevent leakage, harmonize heterogeneous datasets, and reduce dimensionality; (ii) empirical evidence that patient‑level leakage and narrow cohort blocking inflate reported EEG‑based PD detection performance and undermine clinical validity; and (iii) demonstration that learned representations align with canonical EEG frequency bands, yielding interpretable models that achieve 80.6% accuracy across datasets.
@@
Collectively, this framework supports clinical translation by enabling reproducible biomarker validation across diverse cohorts, facilitating integration with multimodal data (e.g., EMG and wearable sensors), and supporting scalable digital health pipelines for monitoring neurodegenerative disease progression in real‑world settings using ML.
@@ II Related Work and Methodological Pitfalls
A growing body of work applies ML techniques to early PD detection using EEG, EMG, and other biosignal modalities, with several studies reporting classification accuracies exceeding 85%, suggesting strong translational potential [17, 18]. However, many of these results rely on evaluation protocols that fail to enforce patient‑level stratification, allowing temporal windows from the same subject to appear in both training and test sets and introducing data leakage. This enables models to exploit subject‑specific idiosyncrasies rather than learning generalizable disease markers, inflating reported performance estimates [15, 19, 20, 21].
Beyond data leakage, EEG‑based PD detection is limited by small sample sizes, accuracy‑only reporting, and preprocessing‑induced bias. Minor subject misallocation can substantially inflate performance [20], while reliance on accuracy obscures clinically relevant trade‑offs [22, 23]. Preprocessing choices further bias results when hyperparameters are tuned on overlapping data [19, 24]. Although nested cross‑validation is standard for ensembling and optimization [19, 25], its use in PD detection would enable stability assessment across patient‑level splits and channel configurations, reducing overconfidence [21]. Thusly, no existing studies evaluate non‑aligned cross‑dataset generalization, leaving cross‑protocol transferability unexplored.
Figure 2: Workflow for EEG channel selection: raw EEG is transformed into spectrograms, channels are ranked via inner‑loop cross‑validation, selected electrodes are evaluated in the outer loop, and window‑level predictions are aggregated into a single patient‑level probability.
Deep learning approaches for EEG‑based PD detection often inherit evaluation weaknesses, including non‑independent subject splits that inflate reported performance [26, 13]. Similar bias arises from non‑stratified cross‑validation and single‑population testing in later studies [16, 27, 28]. Although recent work explores cross‑dataset generalization via generative augmentation with reported accuracies near 85% [29], such methods risk amplifying population‑specific artifacts when synthetic data reproduce dataset biases rather than disease‑relevant structure [30].
Among recent studies, Wu et al. [21] represent a methodological advance by explicitly addressing channel selection in EEG‑based PD detection. Their approach evaluates individual channels using a single train–validation split on the UNM dataset [16], retaining electrodes that exceed a fixed performance threshold before retraining on the selected subset. While this provides insight into channel relevance and improves performance under population‑blocked testing, selection is performed once and remains fixed across runs, without accounting for variability from data partitioning or patient‑level heterogeneity. Moreover, the procedure is tightly coupled to a specific dataset and feature representation, limiting robustness across heterogeneous cohorts.
Our framework embeds channel selection within nested, patient‑stratified cross‑validation, replacing single‑pass, dataset‑specific procedures with stability‑adjusted estimates derived from repeated inner‑fold evaluations. All channels are retained during outer‑loop training to preserve shared spatial representations, while inference is restricted to the most reliable electrodes, yielding robust performance without leakage, singular validations, or dataset‑specific tuning and enabling more stringent, clinically relevant assessments of generalization under heterogeneous conditions.
@@ III-A Preprocessing
EEG Signal Preprocessing Steps
Let $\mathbf{s}_{m}(t)$ denote the raw EEG signal for patient $p_{m}$, sampled at frequency $f_{\text{orig}}$. Signals were resampled to $f_{s}=64~\text{Hz}$ to preserve Beta (13133030 Hz) and Theta (4488 Hz) activity, yielding $\tilde{\mathbf{s}}_{m}(t)$. Each trace was standardized, $\mathbf{z}_{m}(t)=\frac{\tilde{\mathbf{s}}_{m}(t)-\mu_{m}}{\sigma_{m}},$ and normalized to unit amplitude, $\hat{\mathbf{s}}_{m}(t)=\frac{\mathbf{z}_{m}(t)}{\max_{t}|\mathbf{z}_{m}(t)|}.$ Aligned signals $\hat{\mathbf{s}}_{m}(t)$ were stored with subject metadata to ensure traceability and leakage‑free partitioning across outer folds.
To ensure cross‑dataset comparability, channels were harmonized to a common spatial template. Let $\mathcal{C}=\{c_{1},c_{2},\dots,c_{C}\}$ denote the full electrode set, with $|C|$ defined by the Iowa dataset. Channels were grouped into contiguous anatomical regions, $\mathcal{C}=\mathcal{C}_{\text{frontal}}\cup\mathcal{C}_{\text{fronto‑central}}\cup\mathcal{C}_{\text{central}}\cup\cdots\cup\mathcal{C}_{\text{occipital}},$ and ordered within each region according to standard EEG conventions:
@@
Each channel was assigned an index $\pi:\mathcal{C}\rightarrow\{0,1,\dots,C-1\},$ reflecting its left‑to‑right, anterior‑to‑posterior position. Channels absent from a recording were zero‑padded at index $\pi(c)$, ensuring all inputs conform to a consistent spatial structure across heterogeneous datasets.
Cross‑Validation Schema
We employ a patient‑stratified nested cross‑validation design in which the outer loop estimates generalization to unseen patients and the inner loop supports channel selection (Figure 2). Both loops use StratifiedGroupKFold with patient identifiers as grouping variables, ensuring all samples $\mathcal{W}_{m}$ from a patient $p_{m}$ remain within a single fold. This structure prevents patient and temporal leakage, preserves class balance, and maintains unbiased outer‑loop evaluation by nesting channel selection within the inner loop.
Windowing and Feature Extraction
To accommodate variable recording lengths and conditions, each patient’s EEG signal was segmented into channel‑wise spectrograms (Figure 2). With sampling rate $f_{s}=64~\text{Hz}$, signals were windowed using $L_{w}=16{,}384~\text{samples}\;(\approx 256~\text{s}),$ with shorter recordings zero‑padded and longer recordings segmented using hop size $H_{w}=\frac{L_{w}}{2}=8{,}192~\text{samples}.$ Within each window, the short‑time Fourier transform (STFT) was computed. For a discrete‑time signal $\mathbf{w}[n]\in\mathbb{R}^{L_{w}}$,
@@
where $L$ is the analysis window length, $b$ the hop size, $\mathbf{x}[k]$ the window function, $f$ the frequency bin index, and $i$ the time frame index. This yields a complex time–frequency representation indexed by $(i,f)$, with the number of time frames and frequency resolution given by $T=\left\lfloor\frac{L_{w}}{b}\right\rfloor,\quad\Delta f=\frac{f_{s}}{L}.$ Using a discrete FFT implementation with $n_{\text{fft}}=256$ and hop size $h=64$, each channel produces magnitude spectrograms with $F=\frac{n_{\text{fft}}}{2}=128$ frequency bins and $T=\frac{L_{w}}{h}=256$ time frames, stored as $(F\times T\times 1)=(128\times 256\times 1)$ tensors, converted to decibel scale and normalized.
III-B Inner Loop Channel Selection
As shown in Figure 2, channel selection is embedded within the nested cross‑validation framework to address the high dimensionality of EEG recordings with $C\in\{32,64\}$ electrodes. For each outer‑fold training set, per‑channel accuracies $a_{c,f}$ are computed across inner‑folds $f\in\{1,\dots,F\}$ and averaged to obtain a stability‑adjusted score,